### v2

@@ 4 Experiments
This section presents our study’s experimental setup and results, as well as a discussion of their implications. We evaluate the effectiveness of the proposed approach for emotion recognition on two standard audiovisual speech databases and compare it with state-of-the-art methods. Additionally, we perform an ablation study to analyze the impact of various hyperparameters and architectures on the performance of our method.
We present qualitative audio and visual reconstruction results in Fig. 3(a) and Fig. 3(b), respectively. The reconstructions are illustrated for two distinct masking ratios: 50% and 80%. Additional qualitative results are available online (see the webpage address at the end of Section 1). Furthermore, we conducted a study to measure the reconstruction quality at different masking ratios of the audiovisual speech data. This analysis is detailed in Section 4.2 and demonstrates the effectiveness of leveraging multimodality to enhance reconstruction quality.
@@ 4.1.1 Datasets and preprocessing
To pre-train VQ-MAE-AV, we use the VoxCeleb2 dataset [Chung et al., 2018], which offers a broad range of audiovisual speech data from open-source media, with each video featuring a single speaker. We restricted our dataset use to a subset of around 1000 hours of audiovisual speech, encompassing 2170 different speakers. The test set includes about 100 hours of audiovisual speech data with 117 different speakers.
@@
The VQ-VAE-audio and VQ-VAE-visual models are trained on the VoxCeleb2 dataset. The former is trained on short-time Fourier transform (STFT) power spectrograms (${\color[rgb]{1,0,0}\mathbf{x}^{(a)}}$), while the latter is trained on sequences of RGB images (${\color[rgb]{0,0,1}\mathbf{x}^{(v)}}$) captured at 25 fps, cropped and resized to a resolution of $96\times 96$, and aligned using Face-Alignment [Bulat & Tzimiropoulos, 2017]. To compute the STFT, a Hann window of 64 ms (1024 samples at 16 kHz) and a $68$% overlap are used, resulting in a sample rate of 50 Hz, which is twice the sample rate of the visual modality. This leads to sequences of $D=513$ Fourier coefficients.
Emotional audiovisual speech databases
We fine-tune and evaluate the proposed approaches on two emotional audiovisual speech databases.
@@
    RAVDESS [Livingstone & Russo, 2018]: This English database consists of 1440 videos recorded by 24 professional actors and labeled with eight different emotions (neutral, calm, happy, sad, angry, fearful, disgust, surprised).
@@
    CREMA-D [Cao et al., 2014]: It is an English dataset of 7442 videos recorded by 91 actors. Actors spoke from a selection of 12 sentences. The sentences used one of six emotions (anger, disgust, fear, happy, neutral, and sad) with four different intensities (low, medium, high, and unspecified).
We selected these databases as they are commonly used for the emotion recognition task, and their raw data is accessible. We performed 6-fold and 10-fold cross-validation for the RAVDESS and CREAM-D datasets to ensure a fair comparison with previous works. The speaker identities are different in the fine-tuning and evaluation phases.
@@
Figure 4: Quantitative results of the visual reconstruction (a) and audio reconstruction (b). The line represents the mean value across the test dataset, while the shaded area depicts the standard deviation.
@@ 4.1.2 Model architecture
The VQ-VAE-audio (respectively VQ-VAE-visual) architecture is symmetrical concerning the encoder and the decoder, with three 1D (respectively 2D) convolution for the encoder or transposed convolution for the decoder layers and a residual convolution layer. The VQ-VAE models process each frame independently with no time dependency. For each speech power spectrogram frame of size $D=513$, the VQ-VAE-audio encoder compresses it into a discrete latent vector (a column of ${\color[rgb]{1,0,0}\mathbf{x}_{q}^{(a)}}$) of size $D^{\prime}=64$. For each image frame of size $(H=96,W=96,C=3)$, the VQ-VAE-visual encoder compresses it into a discrete latent representation of size $(H^{\prime}=24,W^{\prime}=24)$. The VQ-VAE-audio and the VQ-VAE-visual codebooks contain, respectively, $k_{a}=256$ and $k_{v}=512$ codes of dimension $e_{a}=8$ and $e_{v}=4$. Such a low dimension is chosen to increase the use of the different codes in the codebook [Yu et al., 2021].
Table 1: Accuracy (%) and F1 score (%) results of audiovisual SER. VQ-MAE-AV is pre-trained with both the generative (3.6.1) and contrastive (3.6.2) loss functions using the Query2Emo pooling strategy. The best scores are in bold, and the second-best scores are underlined.
RAVDESS CREMA-D
Method Accuracy F1 score Method Accuracy F1 score
AV-LSTM [Ghaleb et al., 2019] 65.80 - AV-LSTM [Ghaleb et al., 2019] 72.90 -
MuLT [Tsai et al., 2019] 76.60 77.30 MATER [Ghaleb et al., 2020] 67.20 -
CFN-SR [Fu et al., 2021] 75.76 - RAVER [Goncalves & Busso, 2022] 77.30 -
MATER [Ghaleb et al., 2020] 76.30 - MulT Base [Tran & Soleymani, 2022] 68.87 -
AVT [Chumachenko et al., 2022] 79.20 78.20 MulT Large [Tran & Soleymani, 2022] 70.22 -
MDVAE [Sadok et al., 2024] 79.30 80.70 AV-Gating [Ghaleb et al., 2019] 74.00 -
VQ-MAE-AV (ours) 84.80 84.50 VQ-MAE-AV (ours) 80.40 80.00
Table 2: Performance of VQ-MAE-AV using the self-attention fusion strategy for both encoder and decoder, and using the attention pooling strategy for the fine-tuning. ‘Pre-train’ refers to the training of the VQ-MAE-AV for the unmasking task on the VoxCeleb2 database. ‘Freeze’ refers to the freezing of the VQ-MAE-AV encoder.
Method Pre-train Freeze Accuracy (%)
VQ-MAE-AV 29.6
VQ-MAE-AV 70.5
VQ-MAE-AV 81.5
Table 3: Performance of VQ-MAE-AV using the self-attention fusion strategy for both encoder and decoder, fine-tuned on RAVDESS for different encoder depths.
Method Parm. (M) Acc. (%) F1 score (%)
VQ-MAE-AV-6 8.5 75.7 76.0
VQ-MAE-AV-12 13.5 81.5 80.1
VQ-MAE-AV-16 16.8 82.4 82.4
VQ-MAE-AV-20 20.2 81.3 81.3
Table 4: Performance of VQ-MAE-AV using the self-attention fusion strategy for both encoder and decoder, and using the Query2Emo pooling strategy for the fine-tuning. ‘Generative’ corresponds to the loss function in Eq. 3.6.1 and ‘Contrastive’ corresponds to the loss function in Eq. 3.6.2.
Method Contrastive Generative Accuracy (%)
VQ-MAE-AV 75.2
VQ-MAE-AV 84.3
VQ-MAE-AV 84.8
Table 5: Performance of VQ-MAE-AV without contrastive learning for different encoder and decoder architectures, using the attention pooling strategy for the fine-tuning. SAF stands for self-attention fusion and CAF stands for cross-attention fusion.
Method Param. (M) Encoder Decoder Acc. (%)
VQ-MAE-AV 13.5 SAF SAF 81.5
VQ-MAE-AV 25.0 CAF SAF 82.8
VQ-MAE-AV 18.4 SAF CAF 81.8
VQ-MAE-AV 30.0 CAF CAF 83.0
Table 6: Performance of VQ-MAE-AV using the self-attention fusion strategy for both encoder and decoder, without contrastive learning, and using different pooling strategies for the fine-tuning.
Pooling strategy Accuracy (%) F1 score (%)
Mean Pooling 78.1 78.4
Attention Pooling 81.5 80.1
Query2Emo 84.3 84.8
Table 7: Comparison of performance between unimodal VQ-MAE-A-12 or VQ-MAE-V-12 and multimodal VQ-MAE-AV-12 (using SAF for both encoder and decoder and without contrastive learning).
Method Accuracy (%) F1 score (%)
VQ-MAE-A-12 73.2 72.8
VQ-MAE-V-12 74.1 73.9
VQ-MAE-AV-12 81.5 80.1
@@
The VQ-MAE-AV model uses $L=12$ attention blocks in the encoder and $L^{\prime}=4$ in the decoder. Each self-attention layer of a block is divided into four heads. By default, the parameters of the discrete audio and visual tokens ($d$, $h$, and $w$) are set to 4. We set $t_{a}=10$ and $t_{v}=5$ because the sampling rate of ${\color[rgb]{1,0,0}\mathbf{x}^{(a)}}$ is twice the sampling rate of ${\color[rgb]{0,0,1}\mathbf{x}^{(v)}}$. In the ablation study (Section 4.4), we will explore all possible combinations of the VQ-MAE-AV encoder and decoder, according to the two fusion strategies (self-attention fusion and cross-attention fusion). We will also evaluate the impact of the pooling strategy and contrastive learning.
@@ 4.1.3 Training and fine-tuning details
The VQ-MAE-AV is trained using the AdamW optimizer [Loshchilov & Hutter, 2017] with a cosine scheduler to adjust the learning rate, with a 100-epoch warm-up period. The parameters of the optimizer, similar to He et al. [2022], are $\beta_{2}=0.9$, $\beta_{2}=0.95$, and weight_decay$=0.05$. The base learning rate follows the linear scaling rule [Goyal et al., 2017] $lr=(\texttt{base\_lr}=1e-3)\times(\texttt{batchsize}=128)/256$. We distributed the pre-training of VQ-MAE-AV on 4 NVIDIA HGX A100. The training lasted for 160 epochs, and each epoch took approximately 15 minutes.
@@
For the fine-tuning process, we also use the AdamW optimizer [Loshchilov & Hutter, 2017] with a cosine scheduler to adjust the learning rate and with a 40-epoch warm-up period. The parameters of the optimizer are the same as those used for the pre-training. The base learning rate is 1e-4.
@@ 4.2 Audiovisual speech reconstruction quality
In this experiment, we evaluate the reconstruction quality of VQ-MAE-AV when applied to masked audiovisual speech data. The model is fed with a sequence of tokens, some of which have been masked, and it is used to predict the masked tokens, i.e. to reconstruct the complete audiovisual speech sequence from partially observed tokens. We will study the reconstruction performance for different masking ratios.
For this experiment, we compare VQ-MAE-AV to VQ-MAE-A (A for audio) and VQ-MAE-V (V for visual), which are the unimodal versions of VQ-MAE-AV. The average quality performance for the speech and visual modalities is evaluated using the VoxCeleb2 test set. The Peak Signal-to-Noise Ratio (PSNR in dB) is used to assess the quality of the resynthesized visual data, and the Signal-to-Distortion Ratio (SDR in dB) is used to assess the quality of the resynthesized audio data.
@@ 4.3 Audiovisual emotion recognition
Table 1 compares the emotion recognition performance (accuracy and F1 score) of the proposed VQ-MAE-AV model (using the cross-attention fusion strategy for both the encoder and decoder) with the performance of several audiovisual state-of-the-art methods.
@@
The experimental comparison includes LSTM-based methods [Ghaleb et al., 2019] and Transformer-based methods [Tsai et al., 2019, Chumachenko et al., 2022, Goncalves & Busso, 2022]. MATER [Ghaleb et al., 2020] uses distinct Transformer architectures for each modality, merging the embeddings via mean pooling. CFN-SR [Fu et al., 2021] introduces a cross-modality learning approach, employing self-attention and residual structures to model inter- and intra-modality interactions. We also consider the AVT method [Chumachenko et al., 2022], which uses self-attention fusion and modality dropout to address the challenge of one modality being absent. Additionally, RAVER [Goncalves & Busso, 2022] addresses challenges related to modality alignment, temporal information capture, and handling missing features.
@@
The experimental comparison also includes unsupervised audiovisual approaches, such as the multimodal dynamical VAE (MDVAE) [Sadok et al., 2024], which employs a hierarchical latent space to segregate static from dynamic information and modality-common from modality-specific information. Additionally, we compare with MulT [Tran & Soleymani, 2022], a self-supervised approach using the Transformer architecture and pre-trained on the VoxCeleb dataset with a pretext task of reconstructing masked frames (with a masking ratio of 15%).
@@
As can be seen in Table 1, the VQ-MAE-AV model with Query2Emo outperforms the most recent methods overall. VQ-MAE-AV achieves $9.04$%, $5.60$% and $5.50$% better accuracy than the CFN-SR [Fu et al., 2021], AVT [Chumachenko et al., 2022] and MDVAE [Sadok et al., 2024] methods for the RAVDESS dataset, respectively. Regarding the CREMA-D dataset, VQ-MAE-AV achieves $7.50$%, $6.40$%, and $3.10$% better accuracy than the AV-LSTM [Ghaleb et al., 2019], AV-Gating [Ghaleb et al., 2019], and RAVER [Goncalves & Busso, 2022]. Overall, the experimental results demonstrate the effectiveness of the proposed audiovisual self-supervised representation learning technique for SER. This indicates that VQ-MAE-AV learns audiovisual representations that are effectively transferable to the emotion recognition task, resulting in improved performance compared to audiovisual state-of-the-art methods.
4.4 Ablation study and model properties
We conduct a series of experiments to evaluate the impact of various hyperparameters and model designs on the emotion recognition performance of the proposed VQ-MAE-AV model. In the following paragraphs, we will present and discuss the findings of these experiments. All the ablation study is done on RAVDESS.

### v3

@@ 4 Experiments
This section starts by presenting the experimental setup, including the datasets and preprocessing, the model architecture, and the training configuration. Then, we present a first experiment that aims to measure the reconstruction quality of the VQ-MAE-AV model when applied to masked audiovisual speech data. We then evaluate the performance of the proposed approach for emotion recognition using four audiovisual speech datasets and comparing with several state-of-the-art methods. Finally, we present the results of an ablation study that measures the impact of various hyperparameters and architecture choices on the performance of our method.
@@ 4.1.1 Datasets and preprocessing
To pre-train VQ-MAE-AV, we use the VoxCeleb2 dataset (Chung et al., 2018), which offers a broad range of audiovisual speech data from open-source media, with each video featuring a single speaker. We restricted our dataset use to a subset of around 1000 hours of audiovisual speech, encompassing 2170 different speakers. The test set includes about 100 hours of audiovisual speech data with 117 different speakers.Wild2
@@
The VQ-VAE-A and VQ-VAE-V models are trained on the VoxCeleb2 dataset. The former is trained on short-time Fourier transform (STFT) power spectrograms (${\color[rgb]{1,0,0}\mathbf{x}^{(a)}}$), while the latter is trained on sequences of RGB images (${\color[rgb]{0,0,1}\mathbf{x}^{(v)}}$) captured at 25 fps, cropped and resized to a resolution of $96\times 96$, and aligned using Face-Alignment (Bulat and Tzimiropoulos, 2017). To compute the STFT, a Hann window of 64 ms (1024 samples at 16 kHz) and a $68$% overlap are used, resulting in a sample rate of 50 Hz, which is twice the sample rate of the visual modality. This leads to sequences of $D=513$ Fourier coefficients.
Emotional audiovisual speech datasets
We fine-tune and evaluate the proposed approach for audiovisual SER in both controlled and in-the-wild conditions, using four datasets that are presented below.
@@
    RAVDESS (Livingstone and Russo, 2018) is a dataset that consists of 1,440 videos recorded by 24 English-speaking actors in controlled conditions. It is labeled with eight emotions: neutral, calm, happy, sad, angry, fearful, disgust, surprised.
@@
    CREMA-D (Cao et al., 2014) is a dataset that consists of 7,442 videos recorded by 91 English-speaking actors in controlled conditions. Actors spoke from a selection of 12 sentences, using one of six emotions (anger, disgust, fear, happy, neutral, and sad) with four different intensities (low, medium, high, and unspecified).
  • 3.
    DFEW (Jiang et al., 2020) is a dataset containing 16,372 video clips of English-speaking subjects, annotated with seven emotions: neutral, happy, sad, surprise, fear, anger, and disgust. The clips are extracted from diverse movie scenes, ensuring a wide range of variations in expressions, lighting, and occlusions, providing a challenging benchmark for emotion recognition in the wild.
  • 4.
    Aff-Wild2 (Kollias and Zafeiriou, 2018): Aff-Wild2 comprises 564 videos extracted from Youtube (around 2.8M frames), including 554 different English-speaking subjects, and annotated with seven emotions: neutral, anger, disgust, fear, happiness, sadness, surprise. Similar to DFEW, Aff-Wild2 captures spontaneous expressions in uncontrolled environments, including diverse recording conditions, making it suitable for emotion recognition in real-world scenarios.
We selected these datasets as they are commonly used for the emotion recognition task, and the raw audiovisual speech data is accessible. We performed 6-fold and 10-fold cross-validation for the RAVDESS and CREAM-D datasets to ensure a fair comparison with previous works. For DFEW and Aff-Wild2, the train, validation, and test splits are already defined. For all datasets, the speaker identities are different in the training and evaluation sets, allowing us to evaluate SER in a speaker-independent setting.
(a) Qualitative reconstruction results for the audio modality. The spectrogram highlighted in the red box represents the original spectrogram. The two spectrograms on the top right represent the spectrograms masked at 50% and 80%, respectively. The reconstructions using VQ-MAE-AV can be seen directly below these masked spectrograms.
(b) Qualitative reconstruction results for the visual modality. The first sequence shows the original video, followed by the next two sequences representing the masked video with a ratio of 50% and its reconstruction using VQ-MAE-AV. The last two sequences represent the masked video with a ratio of 80% and its reconstruction using VQ-MAE-AV.
Fig. 4: Quantitative results of the audio reconstruction (a) and visual reconstruction (b) using the VQ-MAE-AV model.
@@
Fig. 5: Quantitative results of the visual reconstruction (a) and audio reconstruction (b). The line represents the mean value across the test dataset, while the shaded area depicts the standard deviation.
@@ 4.1.2 Model architecture
The VQ-VAE-A (respectively VQ-VAE-V) architecture is symmetrical concerning the encoder and the decoder, with three 1D (respectively 2D) convolution for the encoder or transposed convolution for the decoder layers and a residual convolution layer. The VQ-VAE models process each frame independently with no time dependency. For each speech power spectrogram frame of size $D=513$, the VQ-VAE-A encoder compresses it into a discrete latent vector (a column of ${\color[rgb]{1,0,0}\mathbf{x}_{q}^{(a)}}$) of size $D^{\prime}=64$. For each image frame of size $(H=96,W=96,C=3)$, the VQ-VAE-V encoder compresses it into a discrete latent representation of size $(H^{\prime}=24,W^{\prime}=24)$. The VQ-VAE-A and the VQ-VAE-V codebooks contain, respectively, $k_{a}=256$ and $k_{v}=512$ codes of dimension $e_{a}=8$ and $e_{v}=4$. Such a low dimension is chosen to increase the use of the different codes in the codebook (Yu et al., 2021).
@@
The VQ-MAE-AV model employs an encoder comprising $L$ attention blocks, where $L$ is set to 12 unless otherwise specified, and a decoder featuring 4 attention blocks. Each self-attention layer of a block is divided into 4 heads. By default, the parameters of the discrete audio and visual tokens ($d$, $h$, and $w$) are set to 4. We set $t_{a}=10$ and $t_{v}=5$ because the sampling rate of ${\color[rgb]{1,0,0}\mathbf{x}^{(a)}}$ is twice the sampling rate of ${\color[rgb]{0,0,1}\mathbf{x}^{(v)}}$. In the ablation study (Section 4.4), we will explore all possible combinations of the VQ-MAE-AV encoder and decoder, according to the two fusion strategies (self-attention fusion and cross-attention fusion). We will also evaluate the impact of the pooling strategy and contrastive learning.
@@ 4.1.3 Training and fine-tuning details
The VQ-MAE-AV is trained using the AdamW optimizer (Loshchilov and Hutter, 2017) with a cosine scheduler to adjust the learning rate, with a 100-epoch warm-up period. The parameters of the optimizer, similar to He et al. (2022), are $\beta_{2}=0.9$, $\beta_{2}=0.95$, and weight_decay$=0.05$. The base learning rate follows the linear scaling rule (Goyal et al., 2017) $lr=(\texttt{base\_lr}=1e-3)\times(\texttt{batchsize}=128)/256$. We distributed the pre-training of VQ-MAE-AV on 4 NVIDIA HGX A100. The training lasted for 160 epochs, and each epoch took approximately 15 minutes.
@@
For the fine-tuning process, we also use the AdamW optimizer (Loshchilov and Hutter, 2017) with a cosine scheduler to adjust the learning rate and with a 40-epoch warm-up period. The parameters of the optimizer are the same as those used for the pre-training. The base learning rate is 1e-4.
@@ 4.2 Audiovisual speech reconstruction quality
In this experiment, we evaluate the reconstruction quality of VQ-MAE-AV when applied to masked audiovisual speech data. The model is fed with a sequence of tokens, some of which have been masked, and it is used to predict the masked tokens, i.e. to reconstruct the complete audiovisual speech sequence from partially observed tokens. We are interested in studying the reconstruction performance for different masking ratios. Qualitative results are presented in Fig. 4(a) and Fig. 4(b) for two masking ratios, 50% and 80%. Additional qualitative results are available online (see the webpage address at the end of Section 1).
In this experiment, we compare VQ-MAE-AV to VQ-MAE-A (A for audio) and VQ-MAE-V (V for visual), which are the unimodal versions of VQ-MAE-AV. The average quality performance for the speech and visual modalities is evaluated using the VoxCeleb2 test set. The Peak Signal-to-Noise Ratio (PSNR in dB) is used to assess the quality of the resynthesized visual data, and the Signal-to-Distortion Ratio (SDR in dB) is used to assess the quality of the resynthesized audio data.
@@ 4.3 Audiovisual emotion recognition
Table 1: Audiovisual emotion recognition results in terms of accuracy (%) and F1 score (%) on the RAVDESS and CREMA-D datasets
. The best scores are in bold, and the second-best scores are underlined. RAVDESS CREMA-D Method Accuracy F1 score Method Accuracy F1 score AV-LSTM (Ghaleb et al., 2019) 65.80 - AV-LSTM (Ghaleb et al., 2019) 72.90 - MuLT (Tsai et al., 2019) 76.60 77.30 MATER (Ghaleb et al., 2020) 67.20 - CFN-SR (Fu et al., 2021) 75.76 - AV-Gating (Ghaleb et al., 2019) 74.00 - MATER (Ghaleb et al., 2020) 76.30 - MulT Base (Tran and Soleymani, 2022) 68.87 - AVT (Chumachenko et al., 2022) 79.20 78.20 MulT Large (Tran and Soleymani, 2022) 70.22 - MDVAE (Sadok et al., 2024) 79.30 80.70 RAVER (Goncalves and Busso, 2022) 77.30 - VQ-MAE-AV (ours) 84.80 84.50 VQ-MAE-AV (ours) 80.40 80.00
Table 2: Accuracy (%) and F1 score (%) results on DFEW and Aff-Wild2. The best scores are in bold, and the second-best scores are underlined.
DFEW Aff-Wild2
Method Accuracy F1 score Method Accuracy F1 score
C3D+LSTM (Zhang et al., 2023) 65.17 - CT-VIPL (Liu et al., 2020) 64.00 33.00
ResNet-18+LSTM (Zhang et al., 2023) 64.32 - Multi-modal ER (Jin et al., 2021) 63.00 40.20
T-MEP (Zhang et al., 2023) 68.85 - CNN-RNN (Antoniadis et al., 2021) 66.80 55.50
VQ-MAE-AV (ours) 70.30 69.80 VQ-MAE-AV (ours) 69.70 68.90
In this section, we evaluate the emotion recognition performance of the proposed VQ-MAE-AV model on 4 different datasets, including 2 recorded in the wild, and we compare the model’s performance with that of 15 state-of-the-art methods. The performance is measured in terms of accuracy and F1 score. For this experimental comparison, we chose the best configuration of the VQ-MAE-AV model, which uses the cross-attention fusion strategy for the encoder and decoder, both generative and contrastive loss functions for self-supervised pre-training on VoxCeleb2, and the Query2Emo strategy for supervised fine-tuning. Other configurations will be investigated in the ablation study of Section 4.4.
4.3.1 Results in controlled conditions
@@
We start by discussing the emotion recognition results on RAVDESS and CREMA-D, the two datasets recorded in controlled lab conditions. The experimental comparison includes LSTM-based methods (Ghaleb et al., 2019) and Transformer-based methods (Tsai et al., 2019; Chumachenko et al., 2022; Goncalves and Busso, 2022). MATER (Ghaleb et al., 2020) uses distinct Transformer architectures for each modality, merging the embeddings via mean pooling. CFN-SR (Fu et al., 2021) introduces a cross-modality learning approach, employing self-attention and residual structures to model inter- and intra-modality interactions. We also consider the AVT method (Chumachenko et al., 2022), which uses self-attention fusion and modality dropout to address the challenge of one modality being absent. Additionally, RAVER (Goncalves and Busso, 2022) addresses challenges related to modality alignment, temporal information capture, and handling missing features.
@@
The experimental comparison also includes the multimodal dynamical VAE (MDVAE) (Sadok et al., 2024), which employs an unsupervised hierarchical latent representation to segregate static from dynamic information and modality-common from modality-specific information. Additionally, we compare with MulT (Tran and Soleymani, 2022), a self-supervised approach using the Transformer architecture and pre-trained on the VoxCeleb dataset with a pretext task of reconstructing masked frames (with a masking ratio of 15%).
@@
As can be seen in Table 1, the VQ-MAE-AV model outperforms the most recent methods overall. VQ-MAE-AV achieves $9.04$%, $5.60$% and $5.50$% better accuracy than the CFN-SR (Fu et al., 2021), AVT (Chumachenko et al., 2022) and MDVAE (Sadok et al., 2024) methods for the RAVDESS dataset, respectively. Regarding the CREMA-D dataset, VQ-MAE-AV achieves $7.50$%, $6.40$%, and $3.10$% better accuracy than the AV-LSTM (Ghaleb et al., 2019), AV-Gating (Ghaleb et al., 2019), and RAVER (Goncalves and Busso, 2022).
4.3.2 Results on in-the-wild datasets
We now discuss the emotion recognition results of the proposed approach in more realistic recording conditions, using the two in-the-wild datasets DFEW and Aff-Wild2. The experimental comparison includes methods based on combinations of convolutional and LSTM networks (Liu et al., 2020; Jin et al., 2021; Zhang et al., 2023; Antoniadis et al., 2021) and a Transformer-based method (Zhang et al., 2023).
As can be seen in Table 2, the VQ-MAE-AV model with Query2Emo achieves superior performance compared to recent methods on the DFEW dataset. Specifically, VQ-MAE-AV outperforms the Transformer-based method T-MEP (Zhang et al., 2023) by $1.45\%$ in accuracy. On the Aff-Wild2 dataset, VQ-MAE-AV achieves a $2.9\%$ improvement in accuracy and a $13.4\%$ boost in F1 score compared to the CNN-RNN model of Antoniadis et al. (2021). These results confirm that the VQ-MAE-AV model also performs well on real audiovisual speech data recorded in uncontrolled environments.
Overall, the experimental results demonstrate the effectiveness of the proposed audiovisual self-supervised representation learning technique for SER. This indicates that VQ-MAE-AV learns audiovisual representations that are effectively transferable to the emotion recognition task, resulting in improved performance compared to state-of-the-art methods.
Table 3: Performance of the baseline VQ-MAE-AV model (first row, in blue), compared to its performance when the encoder is frozen during supervised training on RAVDESS (second row) or when self-supervised pre-training on VoxCeleb2 is discarded (third row).
Pre-training Frozen encoder Accuracy (%)
81.5
70.5
29.6
Table 4: Performance, number of parameters (in millions), and floating-point operations (FLOPs, in billions) of the baseline VQ-MAE-AV model (second row, in blue), compared to variations where the number $L$ of attention blocks in the encoder is modified (other rows).
Encoder depth Param. (M) FLOPs (G) runtime (ms) Acc. (%) F1 score (%)
$L=6$ 8.5 5.9 0.37 75.7 76.0
$L=12$ 13.5 9.4 0.67 81.5 80.1
$L=16$ 16.8 11.7 0.86 82.4 82.4
$L=20$ 20.2 14.1 0.98 81.3 81.3
Table 5: Performance of the VQ-MAE-AV model when using the contrastive (1st row), the generative (2nd row), or both (3rd row) loss functions during the self-supervised pre-training.
Contrastive Generative Accuracy (%)
75.2
84.3
84.8
Table 6: Performance and number of parameters of the baseline VQ-MAE-AV model (first row, in blue), compared to variations with different modality-fusion strategies at the encoder and decoder (other rows).
Encoder Decoder Param. (M) Acc. (%)
SAF SAF 13.5 81.5
CAF SAF 25.0 82.8
SAF CAF 18.4 81.8
CAF CAF 30.0 83.0
Table 7: Performance of the baseline VQ-MAE-AV model (1st row, in blue), compared to variations of the emotion recognition model (other rows).
Emotion recognition model Accuracy (%) F1 score (%)
Attention Pooling 81.5 80.1
Mean Pooling 78.1 78.4
Query2Emo 84.3 84.8
Table 8: Performance of the baseline VQ-MAE-AV model using the audio and visual modalities (1st row, in blue), compared to the equivalent model using only the audio modality (2nd row) or the visual modality (3rd row).
Modality Accuracy (%) F1 score (%)
Audio + visual 81.5 80.1
Audio 73.2 72.8
Visual 74.1 73.9
4.4 Ablation study
In this section, we present a series of ablation experiments using the RAVDESS dataset. Our objective is to evaluate the impact of various hyperparameters and model designs on the emotion recognition performance of the proposed VQ-MAE-AV model. In the ablation experiments, we systematically vary one single hyperparameter or model block at a time, starting from a baseline configuration. This baseline uses self-attention fusion for both the encoder and the decoder (see Section 3.5), it uses only the generative cross-entropy loss function for self-supervised pre-training (see Section 3.6), and it relies on attention pooling for supervised fine-tuning (see Section 3.7). This configuration differs from the best-performing one used for the audiovisual emotion recognition experiments presented in Section 4.3, which used the cross-attention fusion strategy for the encoder and decoder, both generative and contrastive loss functions for self-supervised pre-training, and the Query2Emo strategy for supervised fine-tuning. For clarity and ease of interpretation, the baseline configuration is highlighted in blue in the tables of results of the present section.